OpenStudy (anonymous):

solve for x: √(x+12)=x

OpenStudy (anonymous):

@jim_thompson5910 please help, how do I solve?

OpenStudy (anonymous):

@Luis_Rivera

OpenStudy (anonymous):

I got 4 for my answer, is that correct?

OpenStudy (anonymous):

@e.cociuba

OpenStudy (phi):

did you square both sides of the equation ?

OpenStudy (phi):

you would get x+12= x^2 you can re-arrange that to look like x^2 -x -12 =0 now factor to get 2 possible solutions.

OpenStudy (anonymous):

yea, and somehow I got 4, would that be correct?

OpenStudy (phi):

to factor, look at the last number , the -12 list all pairs of factors of 12 1,12 2,6 3,4 the - in front of the 12 means the factors will have different signs the middle number -x or -1*x means the factors add up to -1 of the 3 pair: +3-4 (the last pair) work you get (x+3)(x-4)=0 this means x= -3 or x= 4

OpenStudy (anonymous):

so if you were to look at my answer choices above 4 is right, correct?

OpenStudy (phi):

but in these equations, people only take the positive square root in other words sqrt(-3+12)= -3 is not accepted. that leaves x=4 sqrt(4+12) is sqrt(16) , and sqrt(16) is 4 4=4 so x=4 is the answer